OpenStudy (anonymous):

Solve by elimination: 3x+4y=1 5x-3y=21 please help.

OpenStudy (anonymous):

you can eliminate x, or y which one do you want to eliminate?

OpenStudy (anonymous):

ok lets eliminate y

OpenStudy (anonymous):

to be able to eliminate the variables must be in the same magnitude

OpenStudy (anonymous):

4y 3y so, we need the least common multiple

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

Or 1

OpenStudy (anonymous):

4 and 3 what is it called .... both can get a number when multiplied we want the smallest one. which one is it ?

OpenStudy (anonymous):

It's 1

OpenStudy (anonymous):

we need as if it was LCD least common denominator

OpenStudy (anonymous):

same problem

OpenStudy (anonymous):

from 3 and 4

OpenStudy (anonymous):

so we can multiply both terms with any number and then they must be the same

OpenStudy (anonymous):

3x+4y=1 5x-3y=21

OpenStudy (anonymous):

So 1 wouldn't be the least Lcd?

OpenStudy (anonymous):

no, 3 and 4 are different!

OpenStudy (anonymous):

3*1 is not equal to 4*1

OpenStudy (anonymous):

This is where I get confused Ohhhh I see what you're saying now

OpenStudy (anonymous):

yeah, I don't know what its called!!! but it works like here\[\frac{ 1 }{ 3 } ,\frac{ 1}{ 4 }\]

OpenStudy (anonymous):

Ohh I see now, now it's making sense lol.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so the smallest common one is 12 I believe

OpenStudy (anonymous):

the 3 fraction needs to be extended by 4, and the 4 fraction by 3.

OpenStudy (anonymous):

we have equations here, not fractions, but it works very similar: 3x+4y=1 5x-3y=21 the above one has 4y. we need to extend the equation by 3 the below one has -3y, so a 3 term. we need to extend eq by 4

OpenStudy (anonymous):

we do this, so that the y terms are then the same type of y term (same number with y)

OpenStudy (anonymous):

9x+12y=3 <- was multiplied by 3 20x-12y=84 <- was multiplied by 4

OpenStudy (anonymous):

Okay, A bit complicated but I'm getting it, there we go. I was about to ask where do we multiply those terms

OpenStudy (anonymous):

But that's not it right?

OpenStudy (anonymous):

now we add them together :) that's the "elimination " part: 9x+12y=3 + 20x-12y=84 ------------- 29 x = 87 the y was actually "eliminated" when adding, because it was once positive once negative, with same magnitude

OpenStudy (anonymous):

This make's sense. It's like how my teacher did it but I didn't know it was being eliminated when adding

OpenStudy (anonymous):

So how do we get the x, and y from here?

OpenStudy (anonymous):

we have 29 x 29 x = 87 well we want x, so we should divide both sides by 29

OpenStudy (anonymous):

which gives us 3

OpenStudy (anonymous):

yep and then, you can plug that in any of the two original equations and solve for y, to also have that

OpenStudy (anonymous):

Thanks so much, I really needed this

OpenStudy (anonymous):

glad we figured it out :)

OpenStudy (anonymous):

Wait, now one sec lol. So when I solve for y it's the exact same thing?

OpenStudy (anonymous):

He left me lol, I'm still a bit confused.

OpenStudy (anonymous):

Please help? lol .

OpenStudy (anonymous):

3x + 4y = 1 ------- [1] 5x - 3y = 21 -------[2] [1] x 5 ---- 15x + 20y = 5 ---- [3] [2] x 3 ---- 15x - 9y = 63 ----- [4] [3] - [4] ----- 20y + 9y = 5 - 63 29y = -58 y = -2 substitute y= -2 in 1 3x + 4y = 1 3x - 8 = 1 3x = 9 x = 3 x= 3 y= -2

OpenStudy (anonymous):

I love you so much. lol.

OpenStudy (anonymous):

i love it as you undertand.... and give a medal..lol